The worldwide number​ (in millions) of a particular phone sold can be approximated by the function f(x)=-871+398lnx, where x=11 corresponds to 2011.
​(a) What was the number of worldwide sales for the particular phone in ​2015?
​(b) If the model continues to be​ accurate, what was the first full year in which the particular phone sales exceed 254 ​million?
PLEASE ANSWER A & B

Answer by Boreal(15235)

f(4-2015)=-871+398 ln 15=206.80 million phones
254=-871+398 lnx. Isolate the ln x
1125=398 ln x
ln x=2.8266. Without rounding, raise that to the e power
x=16.889 or 17 years or 2017
check
-871+398 ln 17=256.61M which is just greater than 254M.